Solve the System of Equations
2y=3x+8
y=2x+3

Answer by jsmallt9(3758)

2y = 3x + 8
y = 2x + 3

Many methods for solving a system like this are taught. With the second equation already "solved for y", the Substitution Method would be a good choice. The second equation ways that y is equal to 2x+3. We can substitute this expression for y in the other equation:
2(2x + 3) = 3x + 8
Note the use of parentheses. It is an extremely good habit to use parentheses when making substitutions. In this case the parentheses help us see that the 2 in front must be distributed:
4x + 6 = 3x + 8
Subtracting 3x:
x + 6 = 8
Subtracting 6:
x = 2

Now we use this value for x and one of the original equations to find the y value:
2y = 3(2) + 8
2y = 6 + 8
2y = 14
Divide by 2:
y = 7

So the solution to this system is the order pair (2, 7)